# Deep Unfolding Network with Spatial Alignment for multi-modal MRI reconstruction

### v1

@@ Abstract
Multi-modal Magnetic Resonance Imaging (MRI) offers complementary diagnostic information, but some modalities are limited by the long scanning time. To accelerate the whole acquisition process, MRI reconstruction of one modality from highly undersampled k-space data with another fully-sampled reference modality is an efficient solution. However, the misalignment between modalities, which is common in clinic practice, can negatively affect reconstruction quality. Existing deep learning-based methods that account for inter-modality misalignment perform better, but still share two main common limitations: (1) The spatial alignment task is not adaptively integrated with the reconstruction process, resulting in insufficient complementarity between the two tasks; (2) the entire framework has weak interpretability. In this paper, we construct a novel Deep Unfolding Network with Spatial Alignment, termed DUN-SA, to appropriately embed the spatial alignment task into the reconstruction process. Concretely, we derive a novel joint alignment-reconstruction model with a specially designed cross-modal spatial alignment term. By relaxing the model into cross-modal spatial alignment and multi-modal reconstruction tasks, we propose an effective algorithm to solve this model alternatively. Then, we unfold the iterative steps of the proposed algorithm and design corresponding network modules to build DUN-SA with interpretability. Through end-to-end training, we effectively compensate for spatial misalignment using only reconstruction loss, and utilize the progressively aligned reference modality to provide inter-modality prior to improve the reconstruction of the target modality. Comprehensive experiments on three real datasets demonstrate that our method exhibits superior reconstruction performance compared to state-of-the-art methods.
@@ 1 Introduction
Magnetic Resonance Imaging (MRI), due to its non-invasive, high resolution, and significant soft tissue contrast, has become a widely used medical imaging technique. However, the MR scan is relatively slow due to the repetitive acquisition of MR signal spatial encoding and hardware limitations. This time-consuming process can lead to discomfort for patients, causing them to move and introducing motion artifacts in the images, which may negatively affect subsequent disease diagnosis. Therefore, accelerating MRI acquisition is of great importance in clinic practice. To accelerate MRI acquisition, a feasible strategy is to reduce the amount of k-space data collected and then reconstruct fully-sampled images by undersampled data.
@@
The Compressed Sensing MRI (CS-MRI) methods enable the accurate reconstruction from undersampled data at sampling rates significantly below those required by the Nyquist sampling theorem (Lustig et al., 2007). They aim at designing some hand-crafted regularizers based on different priors (e.g., structured sparsity (Lai et al., 2016; Yang et al., 2015), non-local sparsity (Qu et al., 2014; Eksioglu, 2016)) of MR images and formulating them into the algorithm optimization to constrain the solution space. Regardless of its theoretical guarantees, it is challenging to handcraft an optimal regularizer. Alternatively, deep learning-based methods have attracted widespread attention in MRI reconstruction due to their accuracy and speed (Wang et al., 2016; Sriram et al., 2020; Zhu et al., 2018). By learning robust feature representations, the deep learning-based methods have achieved impressive reconstruction performance. However, most deep learning-based methods are a black-box process and lack interpretability, which is required in clinical practice. To alleviate this black-box issue, deep unfolding networks (Yang et al., 2020a; Xin et al., 2022; Huang et al., 2023; Zhang et al., 2022; Jiang et al., 2023b, a) have been proposed to incorporate the imaging model and domain knowledge into the network. They unfold the iterations of an optimization algorithm into deep neural networks, thus making the learning process interpretable.
@@
Despite the potential of deep unfolding networks in MRI reconstruction, most of them focus on utilizing information from a single modality. However, in clinical practice, it is common to acquire MR images of different contrasts because each modality reveals distinct tissue and organ characteristics, and the complementary information among modalities contributes to more accurate diagnoses. While different modalities of MR images display different signal types, they are spatially corresponding and depict the same anatomical structures. Researches (Xiang et al., 2018; Feng et al., 2023; Bian et al., 2022; Lei et al., 2023) have shown that reconstruction of one modality (target modality) can be improved by utilizing the information from another modality (reference modality). But, these multi-modal MRI reconstruction methods are all based on the assumption that the images are perfectly aligned, which is rare in practice. The misalignment may negatively affect the reconstruction performance due to insufficiently exploring the correlation of different modalities.
@@
To mitigate the misalignment between modalities, Lai et al. (2017) alternating iteration of the registration and the reconstruction to align the reference modality with the intermediate reconstruction result of the under-sampled target modality. However, the use of conventional iterative optimization is relatively time-consuming. Deep learning-based methods that take inter-modality misalignment into consideration perform better in terms of reconstruction and spatial alignment accuracy, but still share two main common limitations: (1) The spatial alignment is not adaptively integrated with the reconstruction process, resulting in insufficient complementarity between the two tasks; (2) the entire framework has weak interpretability. For example, before reconstruction, Xuan et al. (2022) and Liu et al. (2021) simply align the images of the reference modality with the undersampled images of the target modality, neglecting the negative effects that artifacts in the undersampled images can have on spatial alignment.
@@
To alleviate the aforementioned limitations, we propose a Deep Unfolding Network with Spatial Alignment (DUN-SA) framework in this paper. Specifically, we first derive a novel joint alignment-reconstruction model, in which a cross-modal spatial alignment term is developed to compensate for misalignment between modalities. Subsequently, we relax it into cross-modal spatial alignment and multi-modal reconstruction tasks, and propose an optimization algorithm for these two tasks. Concretely, we use the gradient-based algorithm for spatial alignment and the Half-Quadratic Splitting (HQS) algorithm for reconstruction. By alternately optimizing the two tasks, we solve this model. Further, we unfold the iterative steps of the proposed algorithm and design corresponding network modules. Finally, we propose an end-to-end, deep unfolding network with interpretability. The main contributions of this paper are as follows:
@@
    We propose a novel joint alignment-reconstruction model for multi-modal MRI reconstruction, in which a cross-modal spatial alignment term is developed to compensate for misalignment between modalities. Utilizing gradient-based and HQS techniques, we design an optimization algorithm to alternately solve this model.
@@
    By unfolding the iterative steps of the proposed algorithm and integrating them with specially designed network modules, we construct a deep unfolding network, termed DUN-SA, which exhibits clear interpretability.
@@
    We design the Spatial Alignment Augmented Multi-Modal Learning Block (SAAMMLB) to learn inter-modality prior through aligned images from reference modality, and utilize the Denoising Block (DB) to fully exploit intra-modality prior.
@@
    Through extensive experiments on the fastMRI brain, the IXI dataset, and an in-house dataset, we demonstrate that the proposed DUN-SA outperforms existing state-of-the-art methods in terms of quantitative and qualitative reconstruction results.
@@ 2.1 Single-Modal CS-MRI
The single-modal MRI reconstruction problem is to reconstruct the original image from its partial acquisition. Traditionally, model-based methods usually utilize image prior to improve reconstruction performance. Lai et al. (2016); Yang et al. (2015) use wavelet transform to sparsely represent magnetic resonance images in iterative image reconstructions, capitalizing on sparse prior. Qu et al. (2014); Eksioglu (2016) further sparsify magnetic resonance images by exploiting the similarity of image patches, emphasizing non-local regularization. Ravishankar and Bresler (2010); Zhan et al. (2015) propose dictionary learning for adaptively learning the sparsifying transform (dictionary), and reconstructing the image simultaneously from highly undersampled data.
Deep learning methods, with their powerful ability to represent features, can achieve superior reconstruction quality and acceleration compared to non-deep learning approaches. Wang et al. (2016) utilize CNN architectures to learn the mapping relationship between the MR images obtained from zero-filled and fully-sampled k-space data, significantly improving reconstruction speed while maintaining image quality. Yang et al. (2017); Defazio et al. (2020) not only reconstruct images but also retain more image texture details and reduce more image artifacts through adversarial learning. Contrary to models that reconstruct under-sampled inputs in the image domain, Sriram et al. (2020); Zhu et al. (2018) operate on under-sampled k-space, achieving better reconstruction performance. Shaul et al. (2020); Ran et al. (2020); Zhang et al. (2019); Eo et al. (2018) utilize cross-domain networks for image reconstruction, leveraging information in both image domain and k-space, outperforming single-domain methods. Furthermore, Wang et al. (2020) incorporate the wavelet domain, using information across three domains, further improving reconstruction accuracy.
@@ 2.2 Multi-Modal CS-MRI
In multi-modal MRI reconstruction, a reference modality can guide the reconstruction of the target modality. Model-based methods explore the relationship between modalities based on prior knowledge. Ehrhardt and Betcke (2016) introduce two types of total variation based on position and direction, taking into account the structural prior of MR images. Weizman et al. (2016) acknowledge the differences between modalities and proposes an iterative weighted reconstruction approach. Song et al. (2020) present a method based on coupled dictionary learning for multi-contrast MRI reconstruction, effectively utilizing the structural dependencies between different contrasts. Lai et al. (2017) address the misalignment issue between modalities, proposing a novel MRI image reconstruction method that learns prior from multi-contrast images through a graph wavelet representation, modeling it as a bi-level optimization problem to allow misalignment between these images.
Deep learning-based methods (Sun et al., 2019; Xiang et al., 2019; Dar et al., 2020) combine T1-weighted and T2-weighted images as dual-channel inputs to a deep learning model, improving reconstruction quality by leveraging inter-modality complementary information. (Zhou and Zhou, 2020) integrate multi-modal and dual-domain information, further enhancing reconstruction accuracy. Feng et al. (2023) employ a Transformer structure, using multi-head attention mechanism to deeply capture multi-modal information, offering more global information compared to existing CNN-based methods. Methods that take misalignment between modalities into consideration perform better. Liu et al. (2021) utilize CNN to learn rigid transformation for compensating misalignment between unregistered paired multi-modal MR images, thereby making more efficient use of the reference modality. Xuan et al. (2022) employ a spatial alignment network to compensate misalignment between modalities and incorporate a novel loss function that efficiently trains both the spatial alignment network and the reconstruction network simultaneously.
@@ 2.3 Deep Unfolding Network
Deep Unfolding Networks have achieved impressive results in many medical applications (e.g. dynamic MR imaging (Huang et al., 2021), metal artifact reduction (Wang et al., 2023, 2021), MRI Super-resolution (Yang et al., 2023, 2022; Lei et al., 2023) and CS-MRI (Yang et al., 2020a, b; Lei et al., 2023)), thereby attracting widespread attention recently. By unfolding certain optimization algorithms with network modules, They integrate model-based and learning-based methods well.
In single-modal CS-MRI, for example, by unfolding different optimization algorithms such as Alternating Direction Method of Multipliers (ADMM) (Yang et al., 2020a; Huang et al., 2023; Jiang et al., 2023b), Alternating Iterative Shrinkage-thresholding Algorithm(ISTA) (Zhang et al., 2022), Half-Quadratic Splitting (HQS) (Xin et al., 2022; Jiang et al., 2023a) and combining them with different deep neural networks (e.g., CNNs, S-Nets), deep unfolding networks achieve sparsity in the denoising perspective. This approach not only improves reconstruction quality but also offers interpretability. Additionally, research has shown that incorporating multi-modal information can improve the quality of reconstruction, prompting the development of deep unfolding networks for multi-modal MRI reconstruction. Yang et al. (2020b) capitalize on cross-modal prior and combine channel and spatial attention mechanisms to construct proximal operators, significantly enhancing image reconstruction accuracy. Lei et al. (2023) employ Convolutional Sparse Coding (CSC) techniques for multi-modal MRI modeling. It transfers the common texture information from the reference modality images to the target modality images while avoiding the interference of inconsistent information and integrates deep network modules, achieving superior reconstruction results.
@@ 3 Method
In Section 3, we propose a joint alignment-reconstruction model for multi-modal MRI reconstruction and design an optimization algorithm in Section 3.1. Then, we unfold the iterative steps of the proposed algorithm with corresponding network modules and build DUN-SA in Section 3.2. Finally, we introduce the learnable parameters in DUN-SA and the loss function in Section 3.3.
@@ 3.1.1 The objective functions
Let $x\in\mathbb{R}^{N}$ represent the unknown fully-sampled MRI image, and $k\in\mathbb{C}^{M}$ denote the measured fully-sampled frequency-domain signal, also known as k-space. In the classical compressed sensing problem for MRI, $\tilde{k}\in\mathbb{C}^{m}$ is used to represent the undersampled signal, derived by applying a Fourier transform and a masking operator to the fully-sampled image. Consequently, the relationship between the fully-sampled MRI image and the undersampled k-space signal can be expressed as:
@@
where $F_{m}=MF$ represents masked Fourier transform, $F$ is the Fourier transform operator, $M$ is the masking operator, and $n$ is the measurement noise generally considered as Gaussian noise. Additionally, $m\ll M$, and $\frac{m}{M}$ indicates the compression ratio. $m$ is significantly smaller than $M$, which leads to an underdetermined inverse problem.
@@
where the first term is the fidelity term, ensuring that the reconstructed image $x$ is consistent with the undersampled signal $\tilde{k}$ in k-space; the second term $R(x)$ is the regularization term, enforcing prior constraint like sparsity, and $\eta$ is the balancing factor.
@@
where $x_{\text{ref}}$ represents the fully-sampled image from the reference modality, $\Psi(x,x_{\text{ref}})$ models the correlation between $x$ and $x_{\text{ref}}$, and $\lambda$ is the balancing coefficient.
Furthermore, to compensate for the spatial misalignment between $x$ and $x_{\text{ref}}$, we transform the auxiliary term into a cross-modal spatial alignment term, integrating the spatial alignment task into the reconstruction process, yielding the final energy function:
@@
where $\phi$ represents the displacement field, $\mathcal{T}$ denotes the warp operation, and $\mathcal{T}(x_{\text{ref}},\phi)$ represents the aligned reference modality image. We refer to $\Psi(x,\mathcal{T}(x_{\text{ref}},\phi))$ as the cross-modal spatial alignment term, which models the correlation between target modality and aligned reference modality.

### v2

@@ Abstract
Multi-modal Magnetic Resonance Imaging (MRI) offers complementary diagnostic information, but some modalities are limited by the long scanning time. To accelerate the whole acquisition process, MRI reconstruction of one modality from highly under-sampled k-space data with another fully-sampled reference modality is an efficient solution. However, the misalignment between modalities, which is common in clinic practice, can negatively affect reconstruction quality. Existing deep learning-based methods that account for inter-modality misalignment perform better, but still share two main common limitations: (1) The spatial alignment task is not adaptively integrated with the reconstruction process, resulting in insufficient complementarity between the two tasks; (2) the entire framework has weak interpretability. In this paper, we construct a novel Deep Unfolding Network with Spatial Alignment, termed DUN-SA, to appropriately embed the spatial alignment task into the reconstruction process. Concretely, we derive a novel joint alignment-reconstruction model with a specially designed aligned cross-modal prior term. By relaxing the model into cross-modal spatial alignment and multi-modal reconstruction tasks, we propose an effective algorithm to solve this model alternatively. Then, we unfold the iterative stages of the proposed algorithm and design corresponding network modules to build DUN-SA with interpretability. Through end-to-end training, we effectively compensate for spatial misalignment using only reconstruction loss, and utilize the progressively aligned reference modality to provide inter-modality prior to improve the reconstruction of the target modality. Comprehensive experiments on four real datasets demonstrate that our method exhibits superior reconstruction performance compared to state-of-the-art methods.
@@ 1 Introduction
Magnetic Resonance Imaging (MRI), due to its non-invasive, high resolution, and significant soft tissue contrast, has become a widely used medical imaging technique. However, the MR scan is relatively slow due to the repetitive acquisition of MR signal spatial encoding and hardware limitations. This time-consuming process can lead to discomfort for patients, causing them to move and introducing motion artifacts in the images, which may negatively affect subsequent disease diagnosis. Therefore, accelerating MRI acquisition is of great importance in clinic practice. To accelerate MRI acquisition, a feasible strategy is to reduce the amount of k-space data collected and then reconstruct fully-sampled images by under-sampled data.
@@
The Compressed Sensing MRI (CS-MRI) methods enable the accurate reconstruction from under-sampled data at sampling rates significantly below those required by the Nyquist sampling theorem [20]. They aim at designing some hand-crafted regularizers based on different priors (e.g., structured sparsity [16, 42], non-local sparsity [21, 6]) of MR images and formulating them into the algorithm optimization to constrain the solution space. Regardless of its theoretical guarantees, it is challenging to handcraft an optimal regularizer. Alternatively, deep learning-based methods have attracted widespread attention in MRI reconstruction due to their accuracy and speed [32, 28, 49]. By learning robust feature representations, the deep learning-based methods have achieved impressive reconstruction performance. However, most deep learning-based methods are a black-box process and lack interpretability, which is required in clinical practice. To alleviate this black-box issue, deep unfolding networks [43, 37, 11, 46, 15, 14] have been proposed to incorporate the imaging model and domain knowledge into the network. They unfold the iterations of an optimization algorithm into deep neural networks, thus making the learning process interpretable.
@@
Despite the potential of deep unfolding networks in MRI reconstruction, most of them focus on utilizing information from a single modality. However, in clinical practice, it is common to acquire MR images of different contrasts because each modality reveals distinct tissue and organ characteristics, and the complementary information among modalities contributes to more accurate diagnoses. While different modalities of MR images display different signal types, they are spatially corresponding and depict the same anatomical structures. Researches [35, 8, 2, 18] have shown that reconstruction of one modality (target modality) can be improved by utilizing the information from another modality (reference modality). But, these multi-modal MRI reconstruction methods are all based on the assumption that the images are perfectly aligned, which is rare in practice. The misalignment may negatively affect the reconstruction performance due to insufficiently exploring the correlation of different modalities.
@@
To mitigate the misalignment between modalities, [17] alternating iteration of the registration and the reconstruction to align the reference modality with the intermediate reconstruction result of the under-sampled target modality. However, the use of conventional iterative optimization is relatively time-consuming. Deep learning-based methods that take inter-modality misalignment into consideration perform better in terms of reconstruction and spatial alignment accuracy, but still share two main common limitations: (1) The spatial alignment is not adaptively integrated with the reconstruction process, resulting in insufficient complementarity between the two tasks; (2) the entire framework has weak interpretability. For example, before reconstruction, [38] and [19] simply align the images of the reference modality with the under-sampled images of the target modality, neglecting the negative effects that artifacts in the under-sampled images can have on spatial alignment.
@@
To alleviate the aforementioned limitations, we propose a Deep Unfolding Network with Spatial Alignment (DUN-SA) framework in this paper. Specifically, we first derive a novel joint alignment-reconstruction model, in which an aligned cross-modal prior term is developed to compensate for misalignment between modalities. Subsequently, we relax it into cross-modal spatial alignment and multi-modal reconstruction tasks, and propose an optimization algorithm for these two tasks. Concretely, we use the gradient-based algorithm for spatial alignment and the Half-Quadratic Splitting (HQS) algorithm for reconstruction. By alternately optimizing the two tasks, we solve this model. Further, we unfold the iterative stages of the proposed algorithm and design corresponding network modules. Finally, we propose an end-to-end, deep unfolding network with interpretability. The main contributions of this paper are as follows:
@@
    We propose a novel joint alignment-reconstruction model for multi-modal MRI reconstruction, in which an aligned cross-modal prior term is developed to compensate for misalignment between modalities while learning inter-modality prior. Utilizing gradient-based and HQS techniques, we design an optimization algorithm to alternately solve this model.
@@
    By unfolding the iterative stages of the proposed algorithm and integrating them with specially designed network modules, we construct a deep unfolding network, termed DUN-SA, which exhibits clear interpretability.
@@
    We design the Aligned Inter-modality Prior Learning Block (AIPLB) to learn inter-modality prior through aligned images from reference modality, and utilize the Denoising Block (DB) to fully exploit intra-modality prior.
@@
    Through extensive experiments on the fastMRI dataset, the IXI dataset, the In-house dataset, and the BraTs 2018 dataset, we demonstrate that the proposed DUN-SA outperforms existing state-of-the-art methods in terms of quantitative and qualitative reconstruction results.
@@ 2.1 Single-Modal CS-MRI
The single-modal MRI reconstruction problem is to reconstruct the original image from its partial acquisition. Traditionally, model-based methods usually utilize image prior to improve reconstruction performance. [16, 42] use wavelet transform to sparsely represent magnetic resonance images in iterative image reconstructions, capitalizing on sparse prior. [21, 6] further sparsify magnetic resonance images by exploiting the similarity of image patches, emphasizing non-local regularization. [23, 45] propose dictionary learning for adaptively learning the sparsifying transform (dictionary), and reconstructing the image simultaneously from highly under-sampled data.
Deep learning methods, with their powerful ability to represent features, can achieve superior reconstruction quality and acceleration compared to non-deep learning approaches. [32] utilize CNN architectures to learn the mapping relationship between the MR images obtained from zero-filled and fully-sampled k-space data, significantly improving reconstruction speed while maintaining image quality. [39, 4] not only reconstruct images but also retain more image texture details and reduce more image artifacts through adversarial learning. Contrary to models that reconstruct under-sampled inputs in the image domain, [28, 49] operate on under-sampled k-space, achieving better reconstruction performance. [26, 22, 47, 7] utilize cross-domain networks for image reconstruction, leveraging information in both image domain and k-space, outperforming single-domain methods. Furthermore, [33] incorporate the wavelet domain, using information across three domains, further improving reconstruction accuracy.
@@ 2.2 Multi-Modal CS-MRI
In multi-modal MRI reconstruction, a reference modality can guide the reconstruction of the target modality. Model-based methods explore the relationship between modalities based on prior knowledge. [5] introduce two types of total variation based on position and direction, taking into account the structural prior of MR images. [34] acknowledge the differences between modalities and proposes an iterative weighted reconstruction approach. [27] present a method based on coupled dictionary learning for multi-contrast MRI reconstruction, effectively utilizing the structural dependencies between different contrasts. [17] address the misalignment issue between modalities, proposing a novel MRI image reconstruction method that learns prior from multi-contrast images through a graph wavelet representation, modeling it as a bi-level optimization problem to allow misalignment between these images.
Deep learning-based methods [29, 36, 3] combine T1-weighted and T2-weighted images as dual-channel inputs to a deep learning model, improving reconstruction quality by leveraging inter-modality complementary information. [48] integrate multi-modal and dual-domain information, further enhancing reconstruction accuracy. [8] employ a Transformer structure, using multi-head attention mechanism to deeply capture multi-modal information, offering more global information compared to existing CNN-based methods. Methods that take misalignment between modalities into consideration perform better. [19] utilize CNN to learn rigid transformation for compensating misalignment between unregistered paired multi-modal MR images, thereby making more efficient use of the reference modality. [38] employ a spatial alignment network to compensate misalignment between modalities and incorporate a novel loss function that efficiently trains both the spatial alignment network and the reconstruction network simultaneously.
@@ 2.3 Deep Unfolding Network
Deep Unfolding Networks have achieved impressive results in many medical applications (e.g. dynamic MR imaging [12], metal artifact reduction [31, 30], MRI Super-resolution [40, 41, 18] and CS-MRI [43, 44, 18]), thereby attracting widespread attention recently. By unfolding certain optimization algorithms with network modules, They integrate model-based and learning-based methods well.
In single-modal CS-MRI, for example, by unfolding different optimization algorithms such as Alternating Direction Method of Multipliers (ADMM) [43, 11, 15], Alternating Iterative Shrinkage-thresholding Algorithm(ISTA) [46], Half-Quadratic Splitting (HQS) [37, 14] and combining them with different deep neural networks (e.g., CNNs, ResNets, U-nets), deep unfolding networks achieve sparsity in the denoising perspective. This approach not only improves reconstruction quality but also offers interpretability. Additionally, research has shown that incorporating multi-modal information can improve the quality of reconstruction, prompting the development of deep unfolding networks for multi-modal MRI reconstruction. [44] capitalize on cross-modal prior and combine channel and spatial attention mechanisms to construct proximal operators, significantly enhancing image reconstruction accuracy. [18] employ Convolutional Sparse Coding (CSC) techniques for multi-modal MRI modeling. It transfers the common texture information from the reference modality images to the target modality images while avoiding the interference of inconsistent information and integrates deep network modules, achieving superior reconstruction results.
@@ 3 Method
In Section 3, we propose a joint alignment-reconstruction model for multi-modal MRI reconstruction and design an optimization algorithm in Section 3.1. Then, we unfold the iterative stages of the proposed algorithm with corresponding network modules and build DUN-SA in Section 3.2. Finally, we introduce the details of network parameters and network training in Section 3.3.
@@ 3.1.1 The objective functions
Let $x\in\mathbb{C}^{N}$ represent the fully-sampled MRI image, and $k\in\mathbb{C}^{N}$ denote the measured fully-sampled frequency-domain signal with $N$ being the amount of pixels in the acquired image. Here, $k$ is the Fourier transform of $x$, directly measured in MRI data acquisition. In the classical compressed sensing problem for MRI, $\tilde{k}\in\mathbb{C}^{M}$ is used to represent the under-sampled signal, expressed by applying a Fourier transform and a masking operator to the fully-sampled image. Consequently, the relationship between the fully-sampled MRI image and the under-sampled k-space signal can be formulated as:
@@
where $F_{m}=MF$ represents masked Fourier transform, $F$ is the Fourier transform operator, $M$ is the masking operator, and $n$ is the measurement noise generally considered as Gaussian noise. Additionally, $M\ll N$, and $\frac{M}{N}$ indicates the compression ratio. $M$ is significantly smaller than $N$, which leads to an underdetermined inverse problem.
@@
where the first term is the fidelity term, ensuring that the reconstructed image $x$ is consistent with the under-sampled signal $\tilde{k}$ in k-space; the second term $R(x)$ is the regularization term, enforcing prior constraint like sparsity, and $\eta$ is the balancing factor.
@@
where $x_{\text{ref}}$ represents the fully-sampled image from the reference modality of the same region of interest, $\Psi(x,x_{\text{ref}})$ models the correlation between $x$ and $x_{\text{ref}}$, and $\lambda$ is the balancing coefficient.
Furthermore, to compensate for the spatial misalignment between $x$ and $x_{\text{ref}}$, we transform the auxiliary term into an aligned cross-modal prior term, integrating the spatial alignment task into the reconstruction process, yielding the final energy function:
@@
where $\phi$ represents the displacement field, $\mathcal{T}$ denotes the warp operation, and $\mathcal{T}(x_{\text{ref}},\phi)$ represents the aligned reference modality image. We refer to $\Psi(x,\mathcal{T}(x_{\text{ref}},\phi))$ as the aligned cross-modal prior term, which models the correlation between target modality and aligned reference modality. In order to use standard gradient-based methods to optimize $\phi$, we use a differentiable warp operation [13] to compute $\mathcal{T}(x_{\text{ref}},\phi)$. For each position $p$, we compute a new position $p^{\prime}=p+\phi(p)$ in $x_{\text{ref}}$ resulting in the following expression $\mathcal{T}(x_{\text{ref}},\phi)[p]=x_{\text{ref}}[p+\phi[p]]$. Bilinear interpolation is used to sample from non-integer positions, ensuring the smoothness of the transformation.